OpenStudy (anonymous):

how do you do this (x^2x+7)(x-3)?

OpenStudy (anonymous):

Multiply each term in the (x-3) times each term in the \((x^2+x+7)\). Assuming I have written the expressions correctly. Please verify that.

OpenStudy (anonymous):

Use FOIL

OpenStudy (anonymous):

First Outside Inside Last

OpenStudy (anonymous):

sorry its supposed to look like this (x^2+2x+7)(x-3)

OpenStudy (anonymous):

OK, The same instructions apply. First multiply the x by each term in the trinomial and then multiply the -3 by each term in the trinomial.

OpenStudy (anonymous):

x^2-3x^2+2x+-6+2x+6x+7x-21?

OpenStudy (anonymous):

\(x^2-3x^2+2x+-6+2x+6x+7x-21\) `\(x^2-3x^2+2x+-6+2x+6x+7x-21\)` Not quite right. \(x \times x^2 = x^3\)

OpenStudy (anonymous):

Recheck the rest of your multiplication, I think you may have added some instead?

OpenStudy (anonymous):

x^3-3x^2+2x^2-6x+7x-21

OpenStudy (anonymous):

Yes. Now, combine like terms. \(x^3+2x^2+7x-3x^2-6x-21\)

OpenStudy (anonymous):

x^3-x^2-x-21 @gypsy1274

OpenStudy (anonymous):

When adding numbers with opposite signs, subtract the smaller from the larger and take the sign of the larger. \(-6+7=1\) \(x^3-x^2+x-21\)

OpenStudy (anonymous):

i see what i did is that the awnswer and is this the product?

OpenStudy (anonymous):

Yes. That is the correct answer.

OpenStudy (anonymous):

does it matter if the question was written like this: Find the product (x^2x+7)(x-3)

OpenStudy (anonymous):

Order doesn't matter in multiplication or addition.

OpenStudy (anonymous):

I want to recommend a book to you after you finish your test. It will help you to understand all this much better in preparation for your next class. E-Z Alegbra published by Barron's. It is fun to read.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Your Welcome!

OpenStudy (anonymous):

i just notice though on this page im on it saids this is the awnswer x^3-x^2+x-18 : http://www.jamestown.k12.nd.us/school/highschool/carpenter/ia64.pdf

OpenStudy (anonymous):

I think that is a typo. I don't see how they got -18 when there is only one constant term from the multiplication and it is -21

OpenStudy (anonymous):

ok good thank you